### v1

# A General Theory of Static Response for Markov Jump Processes

Abstract
We consider Markov jump processes on a graph described by a rate matrix that depends on various control parameters. We derive explicit expressions for the static responses of edge currents and steady-state probabilities. We show that they are constrained by the graph topology (i.e. the incidence matrix) by deriving response relations (i.e. linear constraints linking the different responses) and topology-dependent bounds. For unicyclic networks, all scaled current sensitivities are between zero and one and must sum to one. Applying these results to stochastic thermodynamics, we derive explicit expressions for the static response of fundamental currents (which carry the full dissipation) to fundamental thermodynamic forces (which drive the system away from equilibrium).
Introduction.—Nonequilibrium steady state (NESS) of Markov jump processes describe a plethora of phenomena [1] and understanding their response to external perturbations has crucial implications [2] across fields, such as biology [3, 4, 5, 6], nanoelectronics [7, 8], and deep learning [9]. When the Markov jump process is produced by thermal noise, near equilibrium, the response is simple and characterized by the dissipation-fluctuation relation (DFR) [10, 11]. But far-from-equilibrium or for non-thermal processes, the response is significantly more involved; see Refs. [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Recently, for the static response, exact results [24, 25] and tight bounds [26, 27, 28] have been derived assuming Arrhenius-like rates. Moreover, the bounds [24] only hold for local responses, i.e. when the perturbation and the observable are assigned to the same transition.
In this Letter, we build a general static response theory for any Markov jump processes, which describes both local and nonlocal responses for arbitrary parameterizations of the rate matrix. We identify a broad class of parameterizations that produce two types of linear constraints, which we call the Summation Response Relation (SRR) and Cycle Response Relation (CRR). The SRR restricts the responses of the edge flux and probability. The form of such constraints does not depend on the topology of the incidence matrix, but the values of the responses involved strongly depend on it. The CRR limits the sum of local responses by the number of fundamental (Schnakenberg [29]) cycles, which are essential topological characteristics. Moreover, the topology defines which static responses among all combinatorial configurations have universal (remarkably simple) bounds. In unicyclic networks, all responses are bounded; for multicyclic systems, our approach identifies bounded and unbounded responses. In concrete examples considered, the sizes of bounded and unbounded sets are comparable. Finally, for Markov jump processes describing a system in contact with thermal reservoirs (i.e. stochastic thermodynamics), we derive an explicit expression for the static response of fundamental currents to fundamental thermodynamic forces. The former characterize the full dissipation and the latter drive the system out-of-equilibrium.
Setup.—We consider a directed graph $\mathcal{G}$ with $N$ nodes and $N_{e}$ edges and a Markov jump process over the discrete set $\mathcal{S}$ of the $N$ states corresponding to the nodes. Then, the edges $e\in\mathcal{E}$ of $\mathcal{G}$ define possible transitions with the probability rates encoded in the rate matrix $\mathbb{W}/\tau$. In this description, the jump from $n$ to $m$ is the edge (arrow) $+e$ with the source $s(+e)=n$ and the tip $t(e)=m$. For the reverse transition, we have $-e$ with $s(-e)=m$ and $t(-e)=n$. Choosing $\tau=1$ the non-diagonal elements $W_{nm}=W_{e}\geq 0$ become the probabilities per unit of time assigned to the edges $e$. We assume that the matrix $\mathbb{W}$ is irreducible [30] and that all transitions are reversible, i.e. $W_{e}\neq 0$ only if $W_{-e}\neq 0$. With the property of diagonal elements $W_{ii}=-\sum_{j\neq i}W_{ij}$ the described system always exhibits a unique steady-state probability $\bm{\pi}=(\pi_{1},\dots,\pi_{N})^{\intercal}$ that satisfies
$\mathbb{W}\cdot\bm{\pi}$ $=\bm{0}\,,$ (1)
with the normalization $\sum_{i=1}^{N}{\pi_{i}}=1$. We define the transition current along the edge $e$ as $j_{e}=W_{+e}\pi_{s(+e)}-W_{-e}\pi_{s(-e)}$. This definition has a matrix form $\bm{j}=\mathbb{\Gamma}\bm{\pi}$, where the matrix $\mathbb{\Gamma}$ has elements $\Gamma_{ei}=W_{+e}\delta_{is(+e)}-W_{-e}\delta_{is(-e)}$ with the Kronecker symbol $\delta$. In analogy to linear chemical reaction networks [31, 32, 33], the rate matrix can be decomposed as $\mathbb{W}=\mathbb{S}\mathbb{\Gamma}$, where $\mathbb{S}$ is the incidence matrix of the directed graph $\mathcal{G}$ with the elements $S_{ie}=\delta_{is(-e)}-\delta_{is(+e)}$.
We introduce parameterization $\mathbb{W}(\bm{p})$ using the vector $\bm{p}=(\dots,p,\dots)^{\intercal}$ made of $N_{p}$ control parameters $p\in\bm{p}$. Thus, the steady-state condition (1) can be written as
$\mathbb{S}\cdot\bm{j}(\bm{p})=\bm{0}\,,$ (2)
where $\bm{j}(\bm{p})\equiv\bm{j}(\bm{p},\bm{\pi}(\bm{p}))$ depends on $\bm{p}$ both explicitly due to $\mathbb{\Gamma}(\bm{p})$ and implicitly due to $\bm{\pi}(\bm{p})$. The central objects of this work are the static responses of a quantity $q(\bm{p})$ (e.g. a probability $\bm{\pi}$ or a current $\bm{j}$) to perturbations of the parameters $p$, i.e. the elements of the vector $\nabla_{\bm{p}}q(\bm{p})=(\dots,d_{p}q,\dots)^{\intercal}$.
Homogeneous parameterization.—To provide a simple illustration of how a given parameterization can give rise to nontrivial relations amongst static responses, let us consider a parameterization $\bm{h}$ of the rate matrix $\mathbb{W}(\bm{h})$ such that:
$\mathbb{W}(\alpha\bm{h})$ $=\alpha^{k}\mathbb{W}(\bm{h})\,,$ (3)
where $k$ is the positive order of the homogeneous function. The fact that $\mathbb{W}(\alpha\bm{h})\bm{\pi}(\alpha\bm{h})=\alpha^{k}\mathbb{W}(\bm{h})\bm{\pi}(\alpha\bm{h})=0$ implies that $\bm{\pi}(\bm{h})=\bm{\pi}(\alpha\bm{h})$ since the solution of Eq. 1 is unique; $\bm{\pi}(\bm{h})$ is therefore an homogeneous function of order zero of $\bm{h}$ which implies the linear relation (Euler’s theorem for $k=0$)
$\sum_{h\in\bm{h}}hd_{h}\bm{\pi}$ $=0\,.$ (4)
This in turn implies that the current is a homogeneous function $\bm{j}(\alpha\bm{h},\bm{\pi}(\alpha\bm{h}))=\alpha^{k}\bm{j}(\bm{h},\bm{\pi}(\bm{h}))$, which is equivalent to:
$\sum_{h\in\bm{h}}hd_{h}\bm{j}$ $=k\bm{j}\,.$ (5)
Equations 4 and 5 are known in metabolic control analysis as summation theorems; see [34, 35, 36, 37]. In that context, enzyme concentrations play the role of the homogeneous parameters.
Matrix approach to static response.— We now turn to arbitrary parameterizations. Our strategy is to use $\sum_{i=1}^{N}\pi_{i}=1$, and thus $d_{p}\pi_{N}=-\sum_{k=1}^{N-1}d_{p}\pi_{k}$, to arrive at $N-1$ independent equations for others $d_{p}\pi_{k}$ with $k\in\hat{\mathcal{S}}=\mathcal{S}\setminus\{N\}$. To solve this linear problem, we introduce the matrix $\mathbb{K}=\hat{\mathbb{S}}\hat{\mathbb{\Gamma}}$, where $\hat{\mathbb{S}}=[S_{ke}]_{\{k,e\}}$ and $\hat{\mathbb{\Gamma}}=[\Gamma_{ek}-\Gamma_{eN}]_{\{e,k\}}$ are reduced matrices with $k\in\hat{\mathcal{S}}$ and $e\in\mathcal{E}$. In Appendix A, we prove that the matrix $\mathbb{K}$ is invertible [see Eq. 53], and derive the following expressions for the probability and current response matrices:
$\mathbb{R}^{\pi}$ $\equiv[d_{p}\pi_{i}]_{\{i,p\}}=-\begin{pmatrix}\mathbb{K}^{-1}\hat{\mathbb{S}}\mathbb{J}\\ -\sum_{k}(\mathbb{K}^{-1}\hat{\mathbb{S}}\mathbb{J})^{\intercal}_{k}\end{pmatrix}\,,$ (6c)
$\mathbb{R}^{j}$ $\equiv[d_{p}j_{e}]_{\{e,p\}}=\mathbb{P}\mathbb{J}\,.$ (6d)
Here, $\sum_{k}(\mathbb{K}^{-1}\hat{\mathbb{S}}\mathbb{J})^{\intercal}_{k}$ denotes the row that is the sum of all rows of the matrix $\mathbb{K}^{-1}\hat{S}\mathbb{J}$; $\{i,p\}$ (resp. $\{e,p\}$) denote the sets of indexes $i\in\mathcal{S}$ (resp. $e\in\mathcal{E}$) for rows and $p\in\bm{p}$ for columns; $\mathbb{J}\equiv[\partial_{p}j_{e}]_{\{e,p\}}$ is the steady state Jacobian
$\mathbb{J}=[\pi_{s(+e)}\partial_{p}W_{e}(\bm{p})-\pi_{s(-e)}\partial_{p}W_{-e}(\bm{p})]_{\{e,p\}}\;,$ (7)
and the matrix $\mathbb{P}=[P_{ee^{\prime}}]_{\{e,e^{\prime}\}}$ is defined as
$\mathbb{P}$ $\equiv\Big{[}\delta_{ee^{\prime}}-\sum_{x,x^{\prime}\in\hat{\mathcal{S}}}\hat{\Gamma}_{ex}(\mathbb{K}^{-1})_{xx^{\prime}}S_{x^{\prime}e^{\prime}}\Big{]}_{\{e,e^{\prime}\}}\,,$
$\mathbb{P}$ $=\mathbb{I}-\hat{\mathbb{\Gamma}}(\hat{\mathbb{S}}\hat{\mathbb{\Gamma}})^{-1}\hat{\mathbb{S}}\,,$ (8)
with $\mathbb{I}$ denoting the identity matrix. Matrix $\mathbb{P}$ is idempotent [$\mathbb{P}^{2}=\mathbb{P}$] and is known as an oblique projection matrix [38]. Since $\hat{\mathbb{S}}\mathbb{P}=0$, we define $\mathbb{B}$ via
$\mathbb{P}$ $\equiv\Big{[}\sum_{\gamma\in\mathcal{C}}c^{\gamma}_{e}B_{\gamma e^{\prime}}\Big{]}_{\{e,e^{\prime}\}}=\mathbb{C}\mathbb{B}\,,$ (9)
where $\mathbb{C}=(\dots,\bm{c}_{\gamma},\dots)$ is the matrix of the fundamental cycles $\bm{c}^{\gamma}$ defined as the right null vectors of the incidence matrix $\mathbb{S}\bm{c}^{\gamma}=\bm{0}$ (note that $\mathbb{S}$ and $\hat{\mathbb{S}}$ share the same $N_{c}$ cycles). A General Theory of Static Response for Markov Jump Processes is crucial in what follows. In fact, the elements of the matrices $\mathbb{R}^{\pi}$ and $\mathbb{R}^{j}$ cover the responses of all edges to the perturbation of arbitrary parameters. The matrix structures on the r.h.s. of A General Theory of Static Response for Markov Jump Processes encode the information on how those responses are connected to each other.
Response relations.—For a vector $\bm{p}$ such that the matrix $\mathbb{J}$ in Eq. 7 is full row rank ($\text{rk}\mathbb{J}=N_{e}$, i.e. $N_{p}\geq N_{e}$), we can always find the right invertible matrix $\mathbb{J}^{+}$ with $\mathbb{J}\mathbb{J}^{+}=\mathbb{I}$:
$\mathbb{J}^{+}=\mathbb{J}^{\intercal}(\mathbb{J}\mathbb{J}^{\intercal})^{-1}\,.$ (10)
Multiplying both sides of Eqs. 6c and 6d on the vector $\mathbb{J}^{+}\bm{j}$, we arrive at $\mathbb{R}^{\pi}\mathbb{J}^{+}\bm{j}=\bm{0}$ and $\mathbb{R}^{j}\mathbb{J}^{+}\bm{j}=\mathbb{P}\bm{j}=\mathbb{I}\bm{j}$. In coordinate form, these relations give rise to the SRRs:
$\sum_{p\in\bm{p}}\phi_{p}d_{p}\bm{\pi}$ $=0\,,\quad\sum_{p\in\bm{p}}\phi_{p}d_{p}\ln\bm{j}=\bm{1}\,,$ (11)
where $d_{p}\ln\bm{j}=(\dots,j_{e}^{-1}d_{p}j_{e},\dots)^{\intercal}$ is the sensitivity vector, $\bm{1}=(\dots,1,\dots)^{\intercal}$, and $\phi_{p}$ are elements of the vector
$\bm{\phi}$ $=\mathbb{J}^{+}\cdot\bm{j}\,.$ (12)
Equations (11) generalize Eqs. 4 and 5 beyond homogeneous parameters. The structure of Eq. 11 is universal, and the parameterization as well as the topology of the graph are encoded only in the vector Eq. 12, which can be calculated explicitly using Eq. 7. The SRR for the current in Eq. 11 constrains the extremum condition for the edge flux $j_{e}$ with respect to the vector $\bm{p}$: the condition $d_{p}j_{e}=0$ for $p\in\bm{p}$ is satisfied only for $j_{e}=0$ on the r.h.s. of Eq. 11. We stress that many physically motivated parameterisations satisfy Eq. 10.
Edge perturbations.—We now restrict our theory to systems with independent parameters at every edge. In this case, every element $p_{e}$ of the vector $\bm{p}$ is assigned to its own edge $e$, which implies that the matrix $\mathbb{J}(\bm{p})=\text{diag}(\dots,\partial_{p_{e}}j_{e},\dots)$ is diagonal and $\mathbb{J}^{+}=\mathbb{J}^{-1}$. Using Eq. 12, we find
$\phi_{e}=(\partial_{p_{e}}j_{e})^{-1}j_{e}\,.$ (13)
With Eq. 13 at hand, we can find analytic expressions for Eq. 11. Moreover, since the matrix $\mathbb{J}$ is diagonal, we can rewrite Eq. 6d as
$P_{ee^{\prime}}=(\phi_{e^{\prime}}/j_{e^{\prime}})d_{p_{e^{\prime}}}j_{e}\,.$ (14)
Equation 14 shows that the elements $P_{ee^{\prime}}$ are scaled responses. In addition, we define the scaled sensitivity as $\phi_{e^{\prime}}d_{p_{e^{\prime}}}\ln j_{e}=(j_{e^{\prime}}/j_{e})P_{ee^{\prime}}$. Thus, we can characterize the responses and sensitivities from the analysis of the projection matrix $\mathbb{P}$.
For edge parameterization, we call the responses $d_{p_{e^{\prime}}}j_{e}$ and $P_{ee^{\prime}}$ nonlocal if the perturbation edge $e^{\prime}$ does not coincide with the observation edge $e$. In Appendix C we prove that the diagonal elements of $\mathbb{P}$ are bounded:
$0\leq P_{ee}\leq 1\,,$ (15)
which implies the bounds for the local sensitivities [$e^{\prime}=e$ in Eq. 14]:
$0\leq\phi_{e}d_{p_{e}}\ln j_{e}\leq 1\,.$ (16)
Equation 16 holds for an arbitrary edge parameterization that extends the bounds found in Ref. [24]. In fact, the bounds in Ref. [24] assume an Arrhenius-like form of the edge perturbations (special form of $\bm{\phi}$). We will discuss them as in the application section, see Fig. 1.
Using the property of idempotent matrices $\text{tr}~{}\mathbb{P}=\text{rk}~{}\mathbb{P}=N_{c}$, we derive the CRRs:
$\sum_{e}\phi_{e}d_{p_{e}}\ln j_{e}=N_{c}\,,$ (17)
where the l.h.s. is the sum of local sensitivities which are bounded by Eq. 16. Therefore, Eqs. 16 and 17 show that if exactly $N_{c}$ local sensitives are saturated, then the other ones must be zero. We will illustrate Eq. 17 with an example in the application to physics section.
Figure 1: (a): Sketch of unimolecular network. (b): The heights of the purple, blue, green, yellow bins correspond to the scaled responses of edge $1$ to perturbation of $F_{1},\dots,F_{4}$, with $W_{\pm e}=\exp(v_{e}\pm F_{e})$ and $\phi_{e}=\tau_{e}/j_{e}$. The responses calculated for $100$ rate-matrices with random $W_{\pm e}$ homogeneously distributed in $0\leq W_{\pm e}\leq 100$.
Unicyclic networks.— We further restrict our theory to systems with a single cycle. Using Eqs. 2 and 9 for the unicyclic system, we have $P_{ee^{\prime}}=c_{e}B_{\gamma e^{\prime}}$ and $j_{e}=\mathcal{J}c_{e}$ where $\mathcal{J}$ is the flux and $c_{e}$ are the elements of the single cycle $\bm{c}$. Using Eq. 14, we find that if $j_{e},j_{e^{\prime}}\neq 0$ then $j_{e}=j_{e^{\prime}}$ and $\phi_{e^{\prime}}d_{p_{e^{\prime}}}\ln j_{e}=P_{e^{\prime}e^{\prime}}$, which results in the following bounds [see Eq. 15]
$0\leq\phi_{e^{\prime}}d_{p_{e^{\prime}}}\ln j_{e}\leq 1\,,$ (18)
where $e,e^{\prime}\in\mathcal{E}$ cover all combinations of the perturbing $e^{\prime}$ and observing $e$ edges. Equation 18 implies that all the scaled sensitivities are positive and must add up to one due to Eq. 11. Since their individual upper bound coincides with their sum, the saturation of one automatically suppresses the others. Unicyclic systems play an important role in understanding molecular motors and metabolic networks; see, for example, the dynein model in [39] and the biochemical switch in [26]. An example of four-state unicyclic network is shown in Fig. 1a. The histogram in Fig. 1b shows that the scaled responses (color bins) sum up to one for the random rate matrices, as predicted by Eqs. 11 and 18. One also sees that, in general, the responses of $\mathcal{J}$ to edge parameters are shared between all edges, while when one tends to saturate, the other ones are suppressed.
Multicycle systems.—Since elements of $\mathbb{C}$ can always be written using $\{0,1\}$, the parametric dependence of the matrix $\mathbb{P}$ is defined by the $N_{c}N_{e}$ elements of the matrix $\mathbb{B}$ (see Eq. 9), which satisfies $\mathbb{C}\mathbb{B}\mathbb{C}=\mathbb{C}$ due to A General Theory of Static Response for Markov Jump Processes. The matrix $\mathbb{C}$ is the full column rank and $\mathbb{B}\mathbb{C}=\mathbb{I}_{c}$ is the identity matrix of size $N_{c}$. We define $\overline{\mathbb{C}}$ as the invertible submatrix of $\mathbb{C}$ and note that it is always possible to define cycles such that ¯$\overline{\mathbb{C}}=\mathbb{I}_{c}$. Therefore, we can write
$\mathbb{B}\begin{pmatrix}\mathbb{I}_{c}\\ \tilde{\mathbb{C}}\end{pmatrix}=(\overline{\mathbb{B}},\tilde{\mathbb{B}})\begin{pmatrix}\mathbb{I}_{c}\\ \tilde{\mathbb{C}}\end{pmatrix}=\mathbb{I}_{c}\,,\quad\overline{\mathbb{B}}=\mathbb{I}_{c}-\tilde{\mathbb{B}}\tilde{\mathbb{C}}\,,$ (23)
which reduces the number of unknown elements of $\mathbb{P}$ to #$\#\text{var}=N_{c}N_{e}-N_{c}^{2}=N_{c}(N-1)$ elements of $\tilde{\mathbb{B}}$.
[…]
#$\#\text{var}-\#\text{ide}=(N_{c}-1)(N-2)+(N_{e}-1-\#\text{ide})=0\,,$ (24)
and is only possible if #$\#\text{ide}=N_{e}-1$ and $N_{c}=1$ (unicyclic models) or $N=2$ (two states models).
Beyond unicyclic and two-state models, only part of the possible responses are bounded by a linear combination of the local responses. To identify which ones, using Eqs. 9 and 23, we write
$(\mathbb{I}_{c}-\tilde{\mathbb{B}}\tilde{\mathbb{C}})_{ee}$ $=P_{ee}\,,\quad e\leq N_{c}\,,$ (25a)
(~$(\tilde{\mathbb{C}}\tilde{\mathbb{B}})_{ee}$ $=P_{ee}\,,\quad e>N_{c}\,,$ (25b)
[…]
Thermodynamic responses.—The rates can be expressed in terms of their symmetric $v_{e}=v_{-e}=\ln\sqrt{W_{+e}W_{-e}}$ and antisymmetric $w_{e}=\ln\sqrt{W_{+e}/W_{-e}}$ ($w_{-e}=\ln\sqrt{W_{-e}/W_{+e}}$) parts as $W_{+e}=\exp(v_{e}+w_{e})$ and $W_{-e}=\exp(v_{e}-w_{e})$. This decomposition is used in stochastic thermodynamics (i.e. for open systems undergoing transitions caused by thermal reservoirs) because the antisymmetric part defines the energetics of the system [40, 41]:
$\ln\frac{W_{+e}}{W_{-e}}=\sum_{i\in\mathcal{S}}E_{i}S_{ie}+F_{e}\,,\quad F_{e}$ $=\sum_{\alpha\in\mathcal{F}}X_{e\alpha}f_{\alpha}\,.$ (26)
The l.h.s. of Eq. 26 is the entropy change in the reservoirs caused by a transition $e$. It can always be split, in the r.h.s., as a change in the Massieu potential $E_{i}$ of the state $i$ and as a non-conservative contribution $F_{e}=-F_{-e}$. The latter can be expressed, in Eq. 26, as a sum over a subset of reservoirs $\mathcal{F}$ where each term consists of an amount of conserved quantities exchanged with the reservoirs during a transition $e$, $X_{e\alpha}$, multiplied by a conjugated (fundamental [40]) thermodynamic force, $f_{\alpha}$ made of differences of intensive fields of the reservoirs such as inverse temperatures or chemical potentials [40]. Using Eq. 26, we find the responses
$d_{f_{\alpha}}j_{e}$ $=\sum_{\rho\in\mathcal{E}}d_{f_{\alpha}}F_{\rho}d_{F_{\rho}}j_{e}=\sum_{\rho\in\mathcal{E}}X_{\rho\alpha}P_{e\rho}\partial_{F_{\rho}}j_{\rho}\;,$ (27)
which can be obtained analytically for known $\partial_{F_{\rho}}j_{\rho}$. Since the (fundamental [40]) currents exchanged with the different reservoirs are the elements of the vector $\bm{I}=\mathbb{X}^{\intercal}j$, their response to changes in the thermodynamic forces has the following form:
$\mathbb{R}^{I}=[d_{f_{\alpha^{\prime}}}I_{\alpha}]_{\{\alpha,\alpha^{\prime}\}}=\mathbb{X}^{\intercal}\mathbb{P}\mathbb{J}(\bm{F})\mathbb{X}=\mathbb{X}^{\intercal}\mathbb{R}^{j}\mathbb{X}\,,$ (28)
where $\alpha,\alpha^{\prime}\in\mathcal{F}$ and $\mathbb{J}(\bm{F})=\text{diag}(\dots,\partial_{F_{e}}j_{e},\dots)$. Equation 28 shows the linear transformation from the edge responses to the responses of the reservoir currents.
Let us assume now that $w_{e}$ is independent of $v_{e}$. This is relevant, for example, for Arrhenius-like rates [24], as well as for electron transfer rates in CMOS transistors [42] or single electron tunneling rates in the wide-band approximation [43]. Equation 26 shows that $w_{e}$ only depends on the perturbation of the energy and forces, but does not depend on the perturbation of the kinetic parameters. Such kinetic $v_{e}$ and energy (thermodynamic forces) $w_{e}$ perturbations will be constrained by Eq. 11. Calculating the partial derivatives $\partial_{w_{e}}j_{e}=\tau_{e}$ and $\partial_{v_{e}}j_{e}=j_{e}$, we find $\phi_{e}=j_{e}/\tau_{e}$ (resp. $\phi_{e}=1$) for $w_{e}$ (resp. $v_{e}$), where $\tau_{e}=W_{e}\pi_{s(+e)}+W_{-e}\pi_{s(-e)}$ is the edge traffic. Inserting $\phi_{e}$ into Eq. 11, we arrive at the Symmetric and Antisymmetric SRRs
$\sum_{e}\frac{j_{e}}{\tau_{e}}d_{w_{e}}\bm{\pi}$ $=\bm{0}\,,\quad\sum_{e}d_{v_{e}}\bm{\pi}=\bm{0}\,,$ (29a)
$\sum_{e}\frac{j_{e}}{\tau_{e}}d_{w_{e}}\ln\bm{j}$ $=\bm{1}\,,\quad\sum_{e}d_{v_{e}}\ln\bm{j}=\bm{1}\,.$ (29b)
We note that unlike the antisymmetric parameterisation, the symmetric one is homogeneous $\bm{h}=(\dots,v_{e},\dots)^{\intercal}$: the symmetric RSSs in A General Theory of Static Response for Markov Jump Processes coincide with Eqs. 4 and 5.
Figure 2: (a): The graph for double QDD model [40] with $W_{\pm e}=\Gamma_{e}(1+\exp(\pm F_{e}))^{-1}$, where $\Gamma_{e}$ are the electronic tunneling rates. (b): The disks indicate the bounded scaled responses ($20$ out of $36$). (c): Illustration of the CRR for (a), where $N_{c}=3$. The heights of color bins (from purple to yellow) correspond to $\phi_{e}d_{F_{e}}\ln j_{e}$ for $e=1,\dots,6$. (d): Numerical calculations of $(\phi_{e^{\prime}}/j_{e^{\prime}})d_{F_{e^{\prime}}}j_{e}$ for bounded configurations shown in $(b)$. (e): Numerical calculations as in (c), but for unbounded configurations. Rates $W_{\pm e}$ are homogeneously distributed in the range $0\leq W_{\pm e}\leq W_{m}$, where $W_{m}=1000$.
Application.—We illustrate our approach using a Double Quantum Dots (DQD) model [40] with three cycles (see the graph in Fig. 2a). The topology of the DQD is defined by:
$\footnotesize\mathbb{S}=\begin{pmatrix}-1&-1&0&0&-1&0\\ 1&0&-1&0&0&-1\\ 0&1&0&-1&1&0\\ 0&0&1&1&0&1\end{pmatrix}\,,\overline{\mathbb{C}}=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}\,,\tilde{\mathbb{C}}=\begin{pmatrix}-1&0&0\\ -1&-1&0\\ 1&0&-1\end{pmatrix}\,.$ (40)
The topology of the cycles corresponds to #$\#\text{ide}=5$. Solving the independent subsystem of Eqs. (A General Theory of Static Response for Markov Jump Processes), we arrive at
$\tilde{\mathbb{B}}=\begin{pmatrix}-P_{44}&P_{22}-P_{55}-1&P_{66}+P_{33}-1\\ \tilde{B}_{21}&P_{22}-1&\tilde{B}_{23}\\ \tilde{B}_{31}&\tilde{B}_{32}&P_{33}-1\end{pmatrix}\,.$ (44)
Calculating the matrix $\mathbb{P}=\mathbb{C}\big{(}\mathbb{I}-\tilde{\mathbb{B}}\tilde{\mathbb{C}},\tilde{\mathbb{B}}\big{)}$, we find all elements $P_{ee^{\prime}}$ which are the linear combination of diagonal elements and bounded. We illustrated them in Fig. 2(b), where more than half of the responses are bounded as $-1\leq(\phi_{e^{\prime}}/j_{e^{\prime}})d_{p_{e^{\prime}}}j_{e}\leq 1$ for $(e,e^{\prime})$ marked by disks in Fig. 2(b). We note that unlike the always postive local responses, the nonlocal ones can be negative; see numerical simulations in Fig. 2(cd).
Future studies.— Our approach provides powerful tools to identify networks that are highly sensitive or extremely resilient to perturbations. It is also ideally suited to study the responses of enzymatic changes (proofreading, sensing) in chemical reaction networks, in particular in conjunction with recently developed circuit theory [44].
Acknowledgements—This research was supported by Fonds National de la Recherche—FNR, Luxembourg: Project ChemComplex (C21/MS/16356329) and T.A. by Project ThermoElectroChem (C23/MS/18060819).
Appendix A Derivation of A General Theory of Static Response for Markov Jump Processes
The total derivative $d_{p}$ of Eq. 2 on $p\in\bm{p}$ reads
$\mathbb{S}(\partial_{p}\mathbb{\Gamma})\bm{\pi}+\mathbb{S}\mathbb{\Gamma}d_{p}\bm{\pi}=\bm{0}\,.$ (45)
Using $\sum_{i=1}^{N}\pi_{i}=1$ we find one of the derivatives, say the $N$th one, as $d_{p}\pi_{N}=-\sum_{i=1}^{N-1}d_{p}\pi_{i}$. Inserting it in the coordinate form of Eq. 45 we arrive at:
$\footnotesize\sum_{j\in\hat{\mathcal{S}}}\sum_{e\in\mathcal{E}}S_{ie}(\Gamma_{ej}-\Gamma_{eN})d_{p}\pi_{j}$ $=-\sum_{e\in\mathcal{E}}\sum_{k\in\mathcal{S}}S_{ie}(\partial_{p}\Gamma_{ek})\pi_{k}\,,$
$\sum_{j\in\hat{\mathcal{S}}}K_{ij}d_{p}\pi_{j}$ $=-\sum_{k\in\mathcal{S}}S_{ie}J_{ep}\,,$ (46)
where $\hat{\mathcal{S}}=\mathcal{S}\setminus\{N\}$ denotes the space of $N-1$ states, $J_{ep}$ are elements of the Jacobian defined in Eq. 7 and where we introduce the matrix
$\footnotesize\mathbb{K}$ $=\Big{[}\sum_{e\in\mathcal{E}}S_{ie}(\Gamma_{ej}-\Gamma_{eN})\Big{]}_{i,j\in\hat{\mathcal{S}}}=\hat{\mathbb{S}}\hat{\mathbb{\Gamma}}\,,$ (47)
with $\hat{\mathbb{S}}=[S_{ie}]_{i\in\hat{\mathcal{S}},e\in\mathcal{E}}$ and $\hat{\mathbb{\Gamma}}=[\Gamma_{ej}-\Gamma_{eN}]_{e\in\mathcal{E},j\in\hat{\mathcal{S}}}$.
We proceed with the proof that the matrix $\mathbb{K}$ is invertible. The auxiliary matrix $\mathbb{K}_{N}$ coincides with $\mathbb{W}$, except the $N$th row, which is the row of ones. We first prove that $\det\mathbb{K}=\det\mathbb{K}_{N}$ as $\det\mathbb{K}_{N}=$
$\footnotesize\det\begin{pmatrix}[W_{m,n}]_{m\in\tilde{S},n\in\mathcal{S}}\\ \bm{1}^{\intercal}\end{pmatrix}=\det\begin{pmatrix}\mathbb{K}&[W_{mN}]_{m\in\tilde{S}}\\ \bm{0}^{\intercal}&1\end{pmatrix}=\det\mathbb{K}\,.$ (52)
In the second, using the results of Ref. [24] for $\det\mathbb{K}_{N}$ we arrive at:
$\det\mathbb{K}=\det\mathbb{K}_{N}=\prod_{i=1}^{N-1}\lambda_{i}\neq 0\,,$ (53)
where $\lambda_{i}$ are nonzero eigenvalues of the matrix $\mathbb{W}$. Thus, Eq. 53 implies that $\mathbb{K}$ is invertible.
Multiplying Appendix A by $\mathbb{K}^{-1}$, we find the responses $d_{p}\pi_{i}$. Using them to calculate $d_{p}\pi_{N}$ and rewriting the result in matrix form, we arrive at Eq. 6c. Similarly with Eq. 45, we calculate the response $d_{p}\bm{j}$ of the current:
$d_{p}\bm{j}$ $=\partial_{p}\bm{j}+\bm{\Gamma}\partial_{p}\bm{\pi}=\partial_{p}\bm{j}+\Big{[}\sum_{i\in\hat{\mathcal{S}}}(\Gamma_{ei}-\Gamma_{eN})d_{p}\pi_{i}\Big{]}_{e\in\mathcal{E}}$
$=(\mathbb{I}-\hat{\mathbb{\Gamma}}(\hat{\mathbb{S}}\hat{\mathbb{\Gamma}})^{-1}\hat{\mathbb{S}})\partial_{p}\bm{j}=\mathbb{P}\partial_{p}\bm{j}\,,$ (54)
which has the matrix form shown in Eq. 6d.
Figure 3: Model of proofreading [6]. The enzyme E is involved in two complex reactions with wrong $\text{EW}^{*}$, EW and right $\text{ER}^{*}$, ER species. The matrix $\mathbb{C}$ is block diagonal and vectors $\bm{c}^{\gamma}$ are orthogonal.
Appendix B Independent diagonal elements
[…]
$\mathbb{C}=\text{diag}\big{(}\dots,\mathbb{C}_{i},\dots\big{)}\,,$ (55)
where $\mathbb{C}_{i}$ are full column rank matrices and all other elements except those matrices are zero. The examples in the main text have only one block $\mathbb{C}_{1}=\mathbb{C}$. The blocks in Eq. 55 correspond to the cycles without common edges; see Fig. 3. We prove that every diagonal block in Eq. 55 adds one constraint to the diagonal elements, that is, equivalent #$\#\text{ide}=N_{e}-N_{b}$, where $N_{b}$ is the number of blocks $\mathbb{C}_{i}$ (disconnected cycles in $\mathcal{G}$). Consider the block $\mathbb{C}_{i}$ of size $n_{i}\times\gamma_{i}$. We use the decomposition $\mathbb{P}=\mathbb{C}\mathbb{B}$ for block $\mathbb{C}_{i}$ as $\mathbb{P}_{i}=\mathbb{C}_{i}\mathbb{B}_{i}$, where $\mathbb{B}_{i}$ and $\mathbb{P}_{i}$ are diagonal blocks of the matrices $\mathbb{B}$ and $\mathbb{P}$, respectively. In addition, we can use the condition $\mathbb{B}\mathbb{C}=\mathbb{I}$ in block form as $\mathbb{B}_{i}\mathbb{C}_{i}=\mathbb{I}_{\gamma_{i}}$, where $\mathbb{I}_{\gamma_{i}}$ is the identity matrix of size $\gamma_{i}$. Calculating the trace of the block $\mathbb{P}_{i}$ we find
$\text{tr}~{}\mathbb{P}_{i}=\text{tr}~{}\mathbb{C}_{i}\mathbb{B}_{i}=\text{tr}~{}\mathbb{B}_{i}\mathbb{C}_{i}=\text{tr}~{}\mathbb{I}_{\gamma_{i}}=\gamma_{i}\,,$ (56)
which implies the linear constraint on the diagonal elements of $\mathbb{P}$, which consist of the block $\mathbb{P}_{i}$.
Appendix C Supplementary: Proof of Eq. 15
Here we prove that the diagonal elements $P_{ee}$ are bounded as $0\leq P_{ee}\leq 1$. We introduce the projection matrix $\mathbb{\Pi}=\mathbb{I}-\mathbb{P}$ with the diagonal elements defined by
$\Pi_{ee}=\sum_{x,x^{\prime}\in\hat{\mathcal{S}}}\hat{\Gamma}_{ex}(\mathbb{K}^{-1})_{xx^{\prime}}S_{x^{\prime}e}\,.$ (57)
One can see that Eq. 15 is equivalent to $0\leq\mathbb{\Pi}_{ee}\leq 1$. Our proof strategy is the following: first, we prove that all diagonal elements $\Pi_{ee}$ are positive; then we show the upper bound $\Pi_{ee}\leq 1$. To do so for all edges $e$ we distinguish two possible cases: (i) the edge $e$ contains the state $N$ as $s(+e)=N$ or $s(-e)=N$; (ii) the edge $s(+e)=n$ and $s(-e)=m$, where $n,m\neq N$.
(i) Edges with $s(+e)=N$ or $s(-e)=N$.—Let us consider the edge $e$ with $s(+e)=N$ and $s(-e)=n$ which corresponds $S_{Ne}=-1$, $\Gamma_{eN}=W_{+e}$ and $S_{ne}=1$, $\Gamma_{en}=-W_{-e}$. We note that all elements of the row $[\Gamma_{ex}-\Gamma_{eN}]_{x\in\hat{\mathcal{S}}}$ are nonzero due to $\Gamma_{eN}\neq 0$. Then Eq. 57 reads
$\Pi_{ee}=$ $\sum_{x\in\hat{\mathcal{S}}}\hat{\Gamma}_{ex}(\mathbb{K}^{-1})_{xn}S_{ne}$
$=$ $-W_{-e}(\mathbb{K}^{-1})_{nn}-W_{+e}\sum_{x\in\hat{\mathcal{S}}}(\mathbb{K}^{-1})_{xn}$
$=$ $\underbrace{-W_{-e}\frac{M_{nn}(\mathbb{K})}{\det\mathbb{K}}}_{T_{1}}\underbrace{-W_{+e}\sum_{x\in\hat{\mathcal{S}}}\frac{(-1)^{n+x}M_{nx}(\mathbb{K})}{\det\mathbb{K}}}_{T_{2}}\,,$ (58)
where the $M_{ij}(\mathbb{K})$ is the $ij$th minor of the matrix $\mathbb{K}$ and where we use $(\mathbb{K}^{-1})_{ij}=(-1)^{i+j}M_{ji}(\mathbb{K})/\det{\mathbb{K}}$. Using Eq. (A3) we write the elements of $\mathbb{K}$ as the following:
$K_{ij}=\sum_{e^{\prime}\in\mathcal{E}}S_{ie^{\prime}}(\Gamma_{e^{\prime}j}-\Gamma_{e^{\prime}N})\,,$ (59)
and proceed analyzing the dependence of the elements in Eq. 59 on the rates $W_{\pm e}$. Using Eq. 59 for $i,j\in\hat{\mathcal{S}}$ we notice that only one element of $\mathbb{K}$ depends on $\Gamma_{en}=-W_{-e}$ namely
$K_{nn}$ $=S_{ne}(\Gamma_{en}-\Gamma_{eN})+\dots=-W_{-e}+\dots\,,$ (60)
where the symbol $\dots$ denotes the contribution which does not depend on the parameter of interest (in this case $W_{-e}$). Also, Eq. 59 shows that $\Gamma_{eN}=W_{+e}$ contributes to all elements $K_{nj}$ in the $n$th row of the matrix $\mathbb{K}$:
$K_{nj}$ $=-W_{+e}+\dots\,,$ (61)
where $\dots$ denotes the terms which do not depend on $W_{+e}$. Calculating the determinant of $\mathbb{K}$ with respect to the $n$th row, we find
$\footnotesize\det{\mathbb{K}}$ $=\sum_{j}(-1)^{n+j}K_{nj}M_{nj}(\mathbb{K})$ (62)
$=-W_{-e}M_{nn}(\mathbb{K})-W_{+e}\sum_{x\in\hat{\mathcal{S}}}(-1)^{n+x}M_{nx}(\mathbb{K})+Q_{1}\,,$
where $Q_{1}$ does not depend on $W_{\pm e}$. Furthermore, since the $n$th row of $\mathbb{K}$ does not contribute to the minors $M_{nj}(\mathbb{K})$, they do not depend on $W_{\pm e}$.
The sign of the determinant is defined by the size of the system as $\text{sgn}\det{\mathbb{K}}=(-1)^{N-1}$; see Eq. (A5). Since $Q_{1}$ does not depend on $W_{\pm e}$, calculating the sign of Eq. 62 with $W_{\pm e}=0$ we arrive at
$\text{sgn}~{}Q_{1}=\text{sgn}\det\mathbb{K}=(-1)^{N-1}\,.$ (63)
We notice that the sign of the term $T_{1}$ does not depend on $W_{-e}$. We use this property to find the sign of $T_{1}$ considering the limit $W_{-e}\to\infty$:
$\text{sgn}~{}T_{1}=-\text{sgn}\lim_{W_{-e}\to\infty}W_{-e}\frac{M_{nn}(\mathbb{K})}{\det\mathbb{K}}=1\,,$ (64)
where we use $\det\mathbb{K}\to-W_{-e}M_{nn}$ as the rate $W_{-e}$ tends to infinity [see Eq. 62]. Similarly, we find $\text{sgn}~{}T_{2}$ calculating the limit $W_{+e}\to\infty$:
$\text{sgn}~{}T_{2}=-\text{sgn}\lim_{W_{+e}\to\infty}W_{+e}\sum_{x\in\hat{\mathcal{S}}}\frac{(-1)^{n+x}M_{nx}(\mathbb{K})}{\det\mathbb{K}}=1\,.$ (65)
Combining Eqs. 64, 65, and C we arrive at $\Pi_{ee}\geq 0$ for $s(+e)=N$. The case $s(-e)=N$ can be considered in a similar way. Thus, we have $\Pi_{ee}\geq 0$ for the edges of case (i). To find an upper bound, we rewrite Appendix C using Eqs. 62 and 63 as:
$\Pi_{ee}=1-\Bigg{|}\frac{Q_{1}}{\det\mathbb{K}}\Bigg{|}\leq 1\,.$ (66)
Thus, $\Pi_{ee}\geq 0$ and Eq. 66 result in $0\leq\Pi_{ee}\leq 1$ for case (i). In terms of the original matrix $\mathbb{P}$ we have $0\leq P_{ee}\leq 1$.
(ii) Edges with $s(\pm e)\neq N$.—Let us consider the edge $e$ with $s(+e)=n$ and $s(-e)=m$, where $n,m\neq N$. In this case, we have $\Gamma_{eN}=0$ and the row $[\Gamma_{ex}-\Gamma_{eN}]_{x\in\hat{\mathcal{S}}}$ contains only two non-zero elements $\Gamma_{en}=W_{+e}$ and $\Gamma_{em}=-W_{-e}$; and the column $[S_{xe}]_{x\in\hat{\mathcal{S}}}$ has two nonzero elements $S_{ne}=-1$ and $S_{me}=1$. Therefore, the diagonal elements read
$\footnotesize\Pi_{ee}=$ $\sum_{i,j=n,m}\Gamma_{ei}(\mathbb{K}^{-1})_{ij}S_{je}$
$=$ $\underbrace{-W_{+e}\frac{M_{nn}(\mathbb{K})}{\det\mathbb{K}}+(-1)^{n+m}W_{+e}\frac{M_{mn}(\mathbb{K})}{\det\mathbb{K}}}_{T_{3}}$
$\underbrace{+(-1)^{m+n}W_{-e}\frac{M_{nm}(\mathbb{K})}{\det\mathbb{K}}-W_{-e}\frac{M_{mm}(\mathbb{K})}{\det\mathbb{K}}}_{T_{4}}\,.$ (67)
To proceed, we analyze the dependence of the matrix $\mathbb{K}$ on the rates $W_{\pm e}$. We find that two elements of $\mathbb{K}$ depend on $W_{+e}$ namely
$K_{nn}$ $=S_{ne}\Gamma_{en}+\dots=-W_{+e}+\dots\,,$ (68a)
$K_{mn}$ $=S_{me}\Gamma_{en}+\dots=W_{+e}+\dots\,,$ (68b)
and another two elements depend on $W_{-e}$ namely
$K_{nm}$ $=S_{ne}\Gamma_{em}+\dots=W_{-e}+\dots\,,$ (69a)
$K_{mm}$ $=S_{me}\Gamma_{em}+\dots=-W_{-e}+\dots\,.$ (69b)
Therefore, the dependence of the matrix $\mathbb{K}$ on $W_{+e}$ (resp. $W_{-e}$) is encoded in the $n$th column (resp. $m$th column). For this reason, the minors $M_{nn}(\mathbb{K})$ and $M_{mn}(\mathbb{K})$ (resp. $M_{nm}(\mathbb{K})$ and $M_{mm}(\mathbb{K})$) do not depend on $W_{+e}$ (resp. $W_{-e}$). Thus, the signs $\text{sgn}~{}T_{3}$ and $\text{sgn}~{}T_{4}$ can be found by considering the limits $W_{+e}\to\infty$ and $W_{-e}\to\infty$, respectively.
For $T_{3}$ we calculate the $\det{\mathbb{K}}$ in the respect to the $n$th column and use Appendix C:
$\det{\mathbb{K}}$ $=\sum_{i}(-1)^{i+n}K_{in}M_{in}(\mathbb{K})$
$=-W_{+e}M_{nn}(\mathbb{K})+(-1)^{n+m}W_{+e}M_{mn}(\mathbb{K})+\dots\,,$ (70)
where the dots denote the contribution which does not depend on $W_{+e}$. Inserting Appendix C in $T_{3}$ from Appendix C and calculating the sign in the limit $W_{+e}\to\infty$ we arrive at
$\text{sgn}~{}T_{3}=\lim_{W_{+e}\to 0}\frac{-W_{+e}M_{nn}(\mathbb{K})+(-1)^{n+m}W_{+e}M_{mn}(\mathbb{K})}{\det\mathbb{K}}=1\,.$ (71)
For $T_{4}$ we calculate $\det{\mathbb{K}}$ in the respect to the $m$th column and use Appendix C to find
$\det{\mathbb{K}}$ $=\sum_{i}(-1)^{i+n}K_{in}M_{im}(\mathbb{K})$
$=(-1)^{n+m}M_{nm}W_{-e}(\mathbb{K})-W_{-e}M_{mm}(\mathbb{K})+\dots\,,$ (72)
where the dots denote the terms which do not depend on $W_{-e}$. Inserting Appendix C in $T_{4}$ from Appendix C and calculating the sign in the limit $W_{-e}\to\infty$ we arrive at
$\text{sgn}~{}T_{4}=\lim_{W_{-e}\to 0}\frac{(-1)^{n+m}W_{-e}M_{nm}(\mathbb{K})-W_{-e}M_{mm}(\mathbb{K})}{\det\mathbb{K}}=1\,.$ (73)
Using Eqs. 71, 73, and C, we prove that all diagonal elements $\Pi_{ee}\geq 0$ are positive.
Finally, we find an upper bound for $\Pi_{ee}$. We modify the matrix $\mathbb{K}$ adding the $m$th row to the $n$th row. In the result, the new $n$th row does not depend on $W_{\pm e}$; see Appendices C and C. This operation does not change the determinant $\det\mathbb{K}$ that allows to write it using the $m$th row
$\det\mathbb{K}=W_{+e}\alpha+W_{-e}\beta+Q_{2}\,,$ (74)
where $\alpha$, $\beta$ and $Q_{2}$ are quantities which do not depend on $W_{\pm e}$. Equation 74 shows that $\det\mathbb{K}$ is the linear function of $W_{\pm e}$ From the other hand, an explicit dependence of $\det\mathbb{K}$ on $W_{+e}$ (resp. on $W_{-e}$) is described by Appendix C (resp. Appendix C). Thus Eq. 74 must has the following form
$\det\mathbb{K}=$ $W_{+e}(-M_{nn}(\mathbb{K})+(-1)^{n+m}M_{mn}(\mathbb{K}))+$
$+W_{-e}((-1)^{n+m}M_{nm}(\mathbb{K})-M_{mm}(\mathbb{K}))+Q_{2}\,.$ (75)
Since $Q_{2}$ does not depend on $W_{\pm e}$ we have $\text{sgn}\det\mathbb{K}=\text{sgn}~{}Q_{2}$, which gives us:
$\Pi_{ee}=1-\Bigg{|}\frac{Q_{2}}{\det\mathbb{K}}\Bigg{|}\leq 1\,,$ (76)
where we used Appendices C and C. Combining $\Pi_{ee}\geq 0$ and Eq. 76 we have $0\leq\Pi_{ee}\leq 1$ for case (ii). Moving back to the original projection matrix $\mathbb{P}=\mathbb{I}-\mathbb{\Pi}$, we prove Eq. 15.